OpenStudy (anonymous):

-2X/X+2 + X/3 = 4/X+2 HOW?????

OpenStudy (anonymous):

okay, add 2x/x+2 to both sides and post what you get

OpenStudy (anonymous):

X/3 = 4/X+2 + 2X/X+2

OpenStudy (anonymous):

ok good. now factor out 1/x+2 from the RHS

OpenStudy (anonymous):

-6x + (x^2 + 2x) = 12 x^2 - 4x = 12 x^2 - 4x - 12 = 0 x+6 x-2 x = -6 x =2

OpenStudy (anonymous):

multiply by a common denominator of (x+2)*3 to cancel out the fractions btw

OpenStudy (anonymous):

THANKS ^_^